# Pair-breaking as the fundamental limit to persistent-current stabilization in fermionic superfluids

### v1

@@ Abstract
where $N_{v}$ is the number of vortices, $\overline{r}=\frac{1}{N_{v}}\sum_{k}r_{k}^{(\textrm{i})}$ is the mean initial radius, and $\Delta r_{k}=r_{k}^{(\textrm{f})}-r_{k}^{(\textrm{i})}$, $\Delta\theta_{k}=\theta_{k}^{(\textrm{f})}-\theta_{k}^{(\textrm{i})}$, $\Delta t_{k}=t_{k}^{(\textrm{f})}-t_{k}^{(\textrm{i})}$ denote the radial, angular, and temporal displacements between the initial (i) moment of vortex nucleation and its annihilation or the end of the simulation (f). These results correspond to a fixed initial winding number $w_{0}=7$. Impurities significantly impact vortex emission and motion, with trajectories influenced by the size and spacing of the impurities. In the absence of impurities, vortices are symmetrically emitted along the inner ring, forming a periodic pattern that propagates freely. Adding impurities modifies the number and dynamics of vortices. Weak impurities usually deflect vortex paths, while stronger ones induce pinning and trigger the release of additional vortices, visible as color variations in Fig. 3c) for $s_{1}=1.33\xi$. As impurity density increases, the radial mobility exhibits contrasting trends for the two impurity sizes. For larger impurities, both radial and angular mobilities decrease steadily with $N_{d}$, reflecting stronger collective pinning (see $N_{d}=6$ in Fig. 3c). For smaller impurities, $v_{r}$ initially grows (up to $N_{d}=6$), and then starts to decrease. Contrary to conventional expectations, these results show that even when vortices are pinned to impurities (immobile), the current or flow energy remains dissipative, see Fig. 2c). This dissipation arises from pair-breaking processes induced by the impurities. Notably, we find a signature of stabilization effect on the flow energy, for small defects and moderate impurity concentrations ($N_{d}>6$), visible as a plateau in $\delta E_{\mathrm{flow}}$ for $s_{2}=0.67\xi$ and $w_{0}=7$. Interestingly, at $N_{d}=16$ and larger impurity size, vortices exhibit tunneling between impurity sites ($v_{\theta}$ saturates at a non-zero value), signaling a distinct dynamical regime.
Conclusions — We have investigated the stability of persistent currents in fermionic superfluids within the BCS regime using time-dependent density functional theory. Our results demonstrate that the stability of supercurrents is fundamentally limited by a pair-breaking threshold, in sharp contrast to Bose–Einstein condensates, where impurity-induced stabilization is possible Xhani et al. (2025b); Pezzè et al. (2024).
Two distinct channels of energy dissipation emerge in fermionic systems: (i) topological relaxation, associated with winding-number decay and vortex proliferation above the critical winding $w_{c}$, and (ii) microscopic pair-breaking processes, which operate even below this threshold. The latter channel is primarily governed by impurity size, as larger impurities enhance the local suppression of the order parameter and accelerate energy loss without necessarily inducing vortex nucleation. In the vortex-nucleation regime, the size and spacing of impurities strongly influence vortex mobility. Larger impurities reduce vortex mobility, indicating a transition to a pinning regime, while closely spaced impurities allow for vortex hopping between pinning sites. These findings provide a microscopic picture of how disorder and geometry jointly control the interplay between dissipation and stability in fermionic superflows.
@@
The results are directly relevant to astrophysical systems such as neutron stars. Although neutron matter is strongly interacting, its pairing gap and coherence length correspond to a weak-coupling BCS regime Carlson et al. (2012). The impurity spacings explored here are comparable to the expected distances between nuclear clusters in the neutron-star crust Negele and Vautherin (1973); Chamel (2012). Hence, the mechanisms identified (pair-breaking dissipation and impurity-mediated pinning) offer insight into vortex dynamics in neutron-star superfluids and their role in phenomena such as pulsar glitches Chamel and Haensel (2008); Haskell and Melatos (2015); Antonopoulou et al. (2022).
Note added in proof. A recent work Xhani et al. (2025b) has shown increased current stability with impurity density and its dependence on the impurity arrangement with bosonic superfluids.

### v2

@@ Abstract
where $N_{v}$ is the number of vortices, $\overline{r}=\frac{1}{N_{v}}\sum_{k}r_{k}^{(\textrm{i})}$ is the mean initial radius, and $\Delta r_{k}=r_{k}^{(\textrm{f})}-r_{k}^{(\textrm{i})}$, $\Delta\theta_{k}=\theta_{k}^{(\textrm{f})}-\theta_{k}^{(\textrm{i})}$, $\Delta t_{k}=t_{k}^{(\textrm{f})}-t_{k}^{(\textrm{i})}$ denote the radial, angular, and temporal displacements between the initial (i) moment of vortex nucleation and its annihilation or the end of the simulation (f).
Impurities strongly influence both vortex emission and motion, with trajectories determined by defect size and spacing. In clean systems, vortices are symmetrically emitted along the inner edge of the ring, forming a periodic, freely propagating pattern [Fig. 1(a) left panel]. The introduction of impurities breaks this symmetry: for $N_{d}=2$, weak scatterers deflect vortex trajectories, whereas stronger ones induce pinning of some of the vortices and trigger the emission of additional vortices, as visible in Fig. 3(c). As impurity density increases, the radial mobility exhibits opposite trends for the two impurity sizes. For larger defects, both radial and angular mobilities decrease steadily with $N_{d}$, as strong pinning centers restrict the motion of all emitted vortices (see $N_{d}=6$ in Fig. 3(c)). For smaller impurities, however, the radial mobility $v_{r}$ initially rises up to $N_{d}=6$: vortex–impurity scattering deflects trajectories outward, forcing vortices to traverse a broader superfluid region and thus enhancing flow dissipation. Contrary to standard expectations, even pinned vortices exhibit finite flow dissipation, as shown in Fig. 2(c). This effect stems from impurity-induced pair-breaking processes. Interestingly, at moderate impurity densities ($N_{d}>6$), dissipation slows down, particularly for smaller defects—manifested as a plateau in $\delta E_{\mathrm{flow}}$ for $s_{2}=0.67\xi$ and $w_{0}=7$—signaling a partial stabilization of flow energy with increasing $N_{d}$. At higher densities ($N_{d}=16$) and for larger impurities, vortices begin tunneling between defect sites, with the angular mobility $v_{\theta}$ saturating at a finite value. This behavior indicates a qualitatively distinct dynamical regime in which vortices remain locally pinned but acquire collective mobility through inter-site hopping. These observations reveal four distinct mobility regimes—deflected trajectories, individual pinning, collective pinning, and inter-site hopping—absent in clean systems. They represent emergent mechanisms governing vortex dynamics and dissipation in fermionic superfluids.
Conclusions — In conclusion, we demonstrate that impurity size and density precisely control the balance between pair breaking and vortex emission/dynamics—the two primary dissipation channels for persistent currents in BCS-regime fermionic superfluid rings. Time-dependent density functional theory reveals how these parameters tune both mechanisms, the vortex-emission critical winding $w_{c}$ and vortex mobility. Below $w_{c}$, denser or larger impurities enhance pair breaking, causing flow energy decay without winding number loss, unlike clean rings. Above $w_{c}$, the vortex-impurity interaction, defined by the interplay of impurity spacing, size, and coherence length, governs whether vortices are deflected, individually or collectively pinned, or tunnel between impurity sites. These dynamics shift the dissipation mechanism—from pair breaking plus mobile vortices at low $N_{d}$ to dominant pair breaking at higher densities. Impurity size further sets the dependence of $w_{c}$ on $N_{d}$: small impurities yield density-independent critical winding, while large ones progressively enhance it toward the clean-ring pair-breaking limit. Impurity engineering thus provides targeted control of superfluid transport, with implications for ultracold atoms and neutron-star superfluids.
@@
Although neutron matter is strongly interacting, its pairing gap and coherence length correspond to a weak-coupling BCS regime Carlson et al. (2012). The impurity spacings explored here are comparable to the expected distances between nuclear clusters in the neutron-star crust Negele and Vautherin (1973); Chamel (2012). Hence, the mechanisms identified (pair-breaking dissipation and impurity-mediated pinning) offer insight into vortex dynamics in neutron-star superfluids and their role in phenomena such as pulsar glitches Chamel and Haensel (2008); Haskell and Melatos (2015); Antonopoulou et al. (2022).